# Ready Jurist One: Benchmarking Language Agents for Legal Intelligence in Dynamic Environments

### v1

@@ 5 Conclusion
This paper introduces J1-E NVS , an environment designed to simulate agents’ legal interactions in various legal scenarios. We further presents J1-E VAL to assess agents’ ability to operate effectively within dynamic environments. Distinct from prior benchmarks, our framework emphasizes interactivity and encompasses a broad range of realistic legal tasks. Experimental results demonstrate the stability and effectiveness of dynamic environments in J1-E NVS . By evaluating a wide range of existing general-purpose and legal-specific LLMs as legal agents, the study uncovers both notable advancements and persistent challenges faced by current agents. Our framework not only functions as an evaluation system but also serves as a reliable environmental toolkit for data generation or reinforcement learning training.
@@ Appendix A Related Work
Legal AI applies artificial intelligence to enhance or automate various tasks in the legal domain, such as legal information extraction Yao et al. [2022], Shen et al. [2024], judgment prediction Huang et al. [2021], and legal question answering Kien et al. [2020], Zhong et al. [2020]. Early approaches Zhong et al. [2020], Ji et al. [2021] relies on task-specific models and curated datasets for specific legal tasks. With LLMs’ strong generalizability Brown et al. [2020], they have shown notable capabilities in the legal domain. For example, Deng et al. Deng et al. [2023] introduced a structured prompting method to enhance judgment prediction. ChatLaw Cui et al. [2023] and LawLLM Yue et al. [2023, 2024] enhance multitasking capabilities through training on legal knowledge and instructions. Recent explorations Yue et al. [2025b], Chen et al. [2024] have employed LLM-driven agents to simulate legal scenarios. However, a key challenge remains: the lack of a unified framework for evaluating agent performance and guiding future progress.
@@
Existing legal assessments of LLM are predominantly static, single-turn paradigms Fei et al. [2024], Li et al. [2024], Guha et al. [2023], Dai et al. [2023], Yue et al. [2024]. Specifically, Law-Eval Yue et al. [2024] evaluates the LLM using multiple-choice questions from law-related exams. LawBenchFei et al. [2024]and LegalBenchGuha et al. [2023] evaluate various dimensions of legal competence by consolidating existing legal tasks into an LLM-adapted format. Furthermore, LAiW Dai et al. [2023] follows the logic of legal practice of Syllogism by dividing the legal capabilities of LLMs into three levels. Inspired by Bloom’s taxonomy, LexEval Li et al. [2024] proposes a new taxonomy of legal cognitive abilities to organize different tasks. Existing benchmarks adopt a static framework centered on a singular dimension of task competence. Such an approach is inadequate for capturing the dynamic nature of interactions and fails to account for the complex, multi-skill demands of realistic scenarios. Consequently, they are insufficient for the emerging paradigm of agent-based evaluation.
@@
By mimicking human senses and vivid performance, LLMs can be leveraged to power role agents with diverse personalities and environmental contexts, ranging from individuals Shao et al. [2023], Argyle et al. [2023] to demographic groups Zhang et al. [2025], Jiang et al. [2023b]. These agents can coordinate, collaborate, exchange information, or even conduct debates, such as providing information through interaction, showing social intelligence Zhou et al. [2023], Mou et al. [2025]. Based on role agents, some works attempt to build more complex scenarios Stade et al. [2024], Fan et al. [2025]. For example, Generative Agents Park et al. [2023] construct a small town with 25 agents. AI hospital Fan et al. [2025] simulates dynamic medical interactions between a doctor and NPCs, including a patient, an examiner, and a chief physician. The powerful role-representation enabled by LLMs lays the foundation for constructing complex community simulations.
@@ B.1 Legal Sources Processing
We utilize two primary legal sources to construct the role-specific settings.
Chinese Judgment Documents. Judgment documents are official court records of judicial proceedings, outcomes, and factual details regarding involving parties. We collect both civil and criminal judgment documents, and employ an extraction model(GPT-4o) to identify and structure relevant legal elements. For civil judgment documents, we extract the following categories of legal elements.
  • Personal information includes two types. For individual-specific elements, we collect name, gender, ethnicity, birthdate, and residential address. For legal entities(corporations), personal attributes include the corporate name, address, and the identity of the legal representative. These elements are subsequently assigned to the roles of plaintiff and defendant in Complaint & Defence Drafting, as well as Civil Court.
  • Case information captures the procedural details of the case, including plaintiff claims, case details, the defendant’s defence against accusations, and evidences.
@@
  • Procurator information comprises the formal charges and procuratorial opinions, which are utilized to instantiate the procurator agent in criminal courts.
@@ B.2 Personality Theory Construction
Real-world clients exhibit behavioral features according to their occupation and experiences. To ensure legal environments’ better alignment with authentic legal services, we construct non-legal roles characteristics based on the Big Five Personality Traits theory, including five dimensions: extraversion, emotional stability, openness, agreeableness, and conscientiousness. Existing researchYue et al. [2025a], Chen et al. [2024] has demonstrated LLM-driven agents’ capability to follow the preset traits and role-play characters, paving way for our implementation. Specifically, we divide each dimension of Big Five Personality traits theory into three levels(high, medium, and low), and leverage GPT-4o to assign them to non-legal role agents based on their individual information. To be specific, for the non-legal role agent in Knowledge Questioning, we exploit occupations as individual information. For Legal Consultation, we utilize background to ascertain the non-legal role agent’s behavioral style. As for the Complaint Drafting, the case details are employed for behavioral styles construction. For Defence Drafting, we base the behavior modeling on the defendant’s statement of defence. To enhance the distinctness of the agents’ portrayals, we randomly reassign the dimensions with the medium level to three levels with the distribution ratio of 2:1:2. We then map five-dimensional traits into behavioral descriptions that feature role agents’ talking style leveraging GPT-4o to model role agents’ behavior.
@@ B.3 Role Profile Configuration
Type II: Specific Characters are modeled with more granular information to satisfy the increased complexity of environments in Level II, including two types of settings:
@@
  • Case-related details depict the actions and statements of parties involved in the case. For plaintiffs, we assign claims and their evidences. For defendants, we include statement of defence and defendants’ evidences.
@@
These information are assigned to the plaintiff agent(check Figure 16 and Figure 17), and the defendant agent(check Figure 18 and Figure 19).
@@
  • Lawyers attend in civil courts and criminal courts, and are assigned information about their clients(i.e., the plaintiff and the defendant), as well as case-related details, including evidences, defence against accusations, and claims.
@@ Appendix C Data Statistics
As shown in Table 4, we meticulously select 322 quality data in J1-E val , with 160 pieces of legal articles, 93 civil judgment documents, and 69 criminal judgment documents. Legal articles are then utilized for agent configurations in Knowledge Questioning and Legal Consultation, with 98 pieces for Knowledge Questioning and 62 pieces for Legal Consultation. Civil judgment documents are utilized for Complaint & Defence Drafting, as well as Civil Court. Criminal judgment documents are employed for Criminal Court setups. Therefore, we include 508 pieces of data in J1-E val .
@@ D.1 Level-I
  • Non-binary score evaluates other questions based on the original responses of the legal agent. If the ground truth is legal provisions, we measure the score by matching method: 10 if the legal provisions are cited by the legal agent, and 0 if not. For ground truths that involve legal phrases, we leverage GPT-4o to generate a score ranging from 0 to 10. The evaluation prompt is shown in Figure 32.
@@ B IN N BIN B IN N BIN F OR D OC F OR D OC P FS J UD R EA L AW P FS C RI J UD R EA L AW
Table 5: Performances among LLM-driven legal agents on J1-E NVS driven by Qwen3-Instruct-32B.
@@ D.2 Level-II
We evaluate the legal document drafted by the legal agent from two perspective.
(1) Format-following score measures the correspondence of the generated document’s format with the given template. We break down the format into a set of labels $L=[L_{1},L_{2},...,L_{k}]$, which are organized in a specific sequential order $S$. As shown in Figure 29, the complaint format consists of labels of Plaintiff, Defendant, Claims, Facts and Legal Grounds, and Evidences and Sources, Names and Addresses of Witness, following a specific sequential order. As for the defence format, check Figure 30, which includes Defendant, defence arguments, case id, parties and cause of action, Evidences and Sources, names and addresses of witnesses. Through matching, we check whether the generated sequential order $S_{agent}$ follows the given order $S$, and score the accuracy of labels by comparing the generated labels $L_{agent}$ with $L$. The final result is calculated with the following equation.
@@
(2) Document score evaluates the quality of obtained information. To be specific, personal information, including plainiff’s information and the defendant’s information (i.e., name, ethnicity, address, birth date, gender), is measured through exact match(1 or 0). As for the other components(i.e., Claims, Evidences, Statement of Defence, Case Details) are evaluated with prompts in Figure 33.
@@ D.3 Level-III
(1) Procedural-following score provides assessments of completeness of procedural stages, each with several mandatory actions. Only when all mandatory actions are executed is a stage considered complete. Specifically, Civil Court is divided into five stages, i.e., court preparation, court investigation, court debate, court mitigation, and court decision. As illustrated in Figure 34 and Figure 35, we evaluate the court preparation stage, court investigation stage, court debate stage, and the court mitigation stage. On the other hand, as illustrated in Figure 36, Criminal Court is divided into three stages: court preparation, court investigation, and court decision. In addition, we notice that some legal agents are prone to unilateral speech, with all the stages executed within one turn. Such performance is deemed as failure in our settings. Through matching, we calculate the completeness of procedural stages and mandatory actions, and use their average as the $PFS$ score.
(2) Judgment score evaluates the quality of the legal agent’s judgment, which encompasses two parts: the inference part and the judgment part. The inference part contains the court’s inference for the judgments as well as the applied laws. We compare the laws applied by the legal agent with the ground truth to determine the correctness, and leverage GPT-4o to evaluate the inference process. The judgment part in Civil Court includes the agent’s ruling over the case, which is evaluated with GPT-4o. The evaluation prompt could be seen in Figure 33. As for the Criminal Court, the judgment part is divided into three parts, where the criminal charge is scored through matching, and the accuracy of predicted fines and sentences is calculated with log deviation with the following equation. The average of the fine deviation and the sentence deviation is calculated as the final $VER$.
@@
Where: $v$ is the log deviation; $l_{p}$ is the agent predicted result; and $l_{a}$ is the ground truth. Higher $v$ indicates greater deviation and thus worse performance of the legal agent.
@@
For adapted LLMs, we speed up inference using vllm Kwon et al. [2023]. Greedy decoding was used across the evaluations. For models smaller than 32B, we run evaluations using 8 * RTX 4090 GPUs with 24GB memory each. For models larger than 32B, we run evaluations using 8 * A100 GPUs with 80GB memory. For some metrics calculating in evaluation, we use GPT-4o 333gpt-40-2024-11-20. To prevent infinite interaction loops caused by limited model capabilities, we set maximum interaction rounds for each scenario: 15 for knowledge 1 questioning, 10 for legal consultation, 20 for complaint drafting, 15 for defence drafting, 50 for civil, and 35 for criminal court.
@@ F.1 Performance among Various Legal Agents
  • Complaint & Defence Drafting. Table 6 and Table 7 reports additional DOC result in Complaint Drafting and Defence Drafting. Both general and legal-specific models perform relatively well on five sub-metrics of DOC, indicating powerful information collection ability through the multi-turn conversation. However, agents driven by legal-specific or smaller models, such as LawLLM, encounter difficulty in accomplishing the assigned tasks, requiring further training.
  • Civil Court. Table 8 reports additional PFS result in Civil Court. GPT-4o and Claude-3.7 demonstrate remarkable performances in STA and ACT, indicating strong procedural-following ability of these models, while legal-specific models, the inference model, and smaller models are unable to follow civil court procedures.
  • Criminal Court. Table 9 reports additional PFS and DOC result in Criminal Court. GPT-4o, Claude-3.7, and Deepseek-v3 score 100 in UNI. In contrast, legal-specific models fail to navigate the criminal court, resulting lower UNI and poorer performance.
@@ F.2 Analysis of Dialogue Turn Constraints on Simulated Courts Performance
We study the impact of maximum dialogue turn constraint on the performance of legal agent in civil court and criminal court driven by Qwen3-Instruct-32B. For civil court, three groups are introduced, where maximum dialogue turn is set to be 35, 40, and 50. For criminal court, we examine legal agent performance under the maximum turn constraint of 30 and 35 turns. As illustrated in Figure 8, the performance of GPT-4o in civil court demonstrates a positive correlation with the increase in dialogue turn constraint, while Deepseek-v3 and Gemma3-Instruct-12B exhibit a decreasing trend with the increase in maximum dialogue turns. The gap in performance could be attributed to the difference in long-range memory, with GPT-4o effectively retains context and executes procedural operations in detail with prolonged dialogue, while Deepseek-v3 and Gemma3-Instruct-12B suffers to retain early dialogue information, subsequently leading to diminished reasoning and judgment prediction performance. For criminal court, as shown in Figure 9, across all evaluated agents, their performance remains largely stable with the increasing dialogue constraint, as most of cases are concluded within 30 turns. The results demonstrate the value of long-range memory for legal agent performance in complex and multi-turn environment settings, which highlights the necessity for future research on a more robust memory mechanism for legal agents.
@@
Table 10 presents the average number of interactions observed across the six environments. Conversely, agents that fail to complete the task tend to have few turns or directly reach the maximum turn limit. For instance, ChatGLM, InternLM3, and LawLLM achieves only 5.54, 1.05, and 2.45 turns on average in C I . The results indicate that weak interactive abilities prevent the models from sustaining focused, purposeful dialogues, often leading to off-topic, incomplete, or procedurally flawed interactions.
@@
  • Across different models. We select three representative models(i.e., GPT-4o, Deepseek-v3 and Gemma3-Instruct-12B), and re-run each model three times under the same settings of the main experiment in both the CD and DD scenarios. As shown in Table 13, performances of these models remain consistent across three runs, demonstrating the robustness of our framework.

### v2

@@ 5 Conclusion
This paper introduces J1-E NVS , an environment designed to simulate agents’ legal interactions in various legal scenarios. We further present J1-E VAL to assess agents’ ability to operate effectively within dynamic environments. Distinct from prior benchmarks, our framework emphasizes interactivity and encompasses a broad range of realistic legal tasks. Experimental results demonstrate the stability and effectiveness of dynamic environments in J1-E NVS . By evaluating a wide range of existing general-purpose and legal-specific LLMs as legal agents, the study uncovers both notable advances and persistent challenges faced by current agents. Our framework not only functions as an evaluation system but also serves as a reliable environmental toolkit for data generation or reinforcement learning training.
@@ Appendix A Related Work
Legal AI applies artificial intelligence to enhance or automate various tasks in the legal domain, such as legal information extraction Yao et al. [2022], Shen et al. [2024], judgment prediction Huang et al. [2021], and legal question answering Kien et al. [2020], Zhong et al. [2020]. Early approaches Zhong et al. [2020], Ji et al. [2021] rely on task-specific models and curated datasets for specific legal tasks. With LLMs’ strong generalizability Brown et al. [2020], they have shown notable capabilities in the legal domain. For example, Deng et al. Deng et al. [2023] introduced a structured prompting method to enhance judgment prediction. ChatLaw Cui et al. [2023] and LawLLM Yue et al. [2023, 2024] enhance multitasking capabilities through training on legal knowledge and instructions. Recent explorations Yue et al. [2025b], Chen et al. [2024] have employed LLM-driven agents to simulate legal scenarios. However, a key challenge remains: the lack of a unified framework for evaluating agent performance and guiding future progress.
@@
Existing legal assessments of LLM are predominantly static, single-turn paradigms Fei et al. [2024], Li et al. [2024], Guha et al. [2023], Dai et al. [2023], Yue et al. [2024]. Specifically, LegalBenchGuha et al. [2023] and LawBenchFei et al. [2024] evaluate various dimensions of legal competence by consolidating existing legal tasks into an LLM-adapted format. Law-Eval Yue et al. [2024] evaluates the LLM using multiple-choice questions from law-related exams. Inspired by Bloom’s taxonomy, LexEval Li et al. [2024] proposes a new taxonomy of legal cognitive abilities to organize different tasks. Furthermore, LAiW Dai et al. [2023] follows the logic of legal practice of Syllogism by dividing the legal capabilities of LLMs into three levels. Existing benchmarks adopt a static framework centered on a singular dimension of task competence. Such an approach is inadequate for capturing the dynamic nature of interactions and fails to account for the complex, multi-skill demands of realistic scenarios. Consequently, they are insufficient for the emerging paradigm of agent-based evaluation.
@@
By mimicking human senses and vivid performance, LLMs can be leveraged to power role agents with diverse personalities and environmental contexts, ranging from individuals Shao et al. [2023], Argyle et al. [2023] to demographic groups Zhang et al. [2025], Jiang et al. [2023b]. These agents can coordinate, collaborate, exchange information, or even conduct debates, such as providing information through interaction, showing social intelligence Zhou et al. [2023], Mou et al. [2025]. Based on role agents, some works attempts to build more complex scenarios Stade et al. [2024], Fan et al. [2025]. For example, Generative Agents Park et al. [2023] construct a small town with 25 agents. AI hospital Fan et al. [2025] simulates dynamic medical interactions between a doctor and NPCs, including a patient, an examiner, and a chief physician. The powerful role-representation enabled by LLMs lays the foundation for constructing complex community simulations.
@@ B.1 Legal Sources Processing
We utilize two primary legal sources to construct role-specific settings.
Chinese Judgment Documents. Judgment documents are official court records of judicial proceedings, outcomes, and factual details regarding parties involved. We collect both civil and criminal judgment documents, and employ an extraction model (GPT-4o) to identify and structure relevant legal elements. For civil judgment documents, we extract the following categories of legal elements.
  • Personal information includes two types. For individual-specific elements, we collect name, gender, ethnicity, birthdate, and residential address. For legal entities(corporations), personal attributes include the corporate name, address, and identity of the legal representative. These elements are subsequently assigned to the roles of plaintiff and defendant in Complaint & Defence Drafting, as well as the Civil Court.
  • Case information captures the procedural details of the case, including plaintiff claims, case details, the defendant’s defence against accusations, and evidence.
@@
  • Procurator information comprises formal charges and procuratorial opinions, which are utilized to instantiate the procurator agent in criminal courts.
@@ B.2 Personality Theory Construction
Real-world clients exhibit behavioral features according to their occupation and experiences. To ensure legal environments’ better alignment with authentic legal services, we construct non-legal roles characteristics based on the Big Five Personality Traits theory, including five dimensions: extraversion, emotional stability, openness, agreeableness, and conscientiousness. Existing researchYue et al. [2025a], Chen et al. [2024] has demonstrated LLM-driven agents’ capability to follow the preset traits and role-play characters, paving the way for our implementation. Specifically, we divide each dimension of Big Five Personality traits theory into three levels(high, medium, and low), and leverage GPT-4o to assign them to non-legal role agents based on their individual information. To be specific, for the non-legal role agent in Knowledge Questioning, we exploit occupations as individual information. For Legal Consultation, we utilize background to ascertain the non-legal role agent’s behavioral style. As for the Complaint Drafting, the case details are employed for behavioral styles construction. For Defence Drafting, we base the behavior modeling on the defendant’s statement of defence. To enhance the distinctness of the agents’ portrayals, we randomly reassign the dimensions with the medium level to three levels with the distribution ratio of 2:1:2. We then map five-dimensional traits into behavioral descriptions that feature role agents’ talking styles leveraging GPT-4o to model role agents’ behavior.
@@ B.3 Role Profile Configuration
Type II: Specific Characters are modeled with more granular information to satisfy the increased complexity of environments at Level II, including two types of settings:
@@
  • Case-related details depict the actions and statements of the parties involved in the case. For plaintiffs, we assign the claims and their evidence. For defendants, we include the statement of defence and defendants’ evidence.
@@
This information is assigned to the plaintiff agent(check Figure 16 and Figure 17), and the defendant agent(check Figure 18 and Figure 19).
@@
  • Lawyers attend in civil courts and criminal courts, and are assigned information about their clients(i.e., the plaintiff and the defendant), as well as case-related details, including evidence, defence against accusations, and claims.
@@ Appendix C Data Statistics
As shown in Table 4, we meticulously select 322 quality data in J1-E val , with 160 pieces of legal articles, 93 civil judgment documents, and 69 criminal judgment documents. Legal articles are then utilized for agent configurations in Knowledge Questioning and Legal Consultation, with 98 pieces for Knowledge Questioning and 62 pieces for Legal Consultation. Civil judgment documents are utilized for Complaint & Defence Drafting, as well as the Civil Court. Criminal judgment documents are employed for Criminal Court setups. Therefore, we include 508 pieces of data in J1-E val .
@@ D.1 Level-I
  • Non-binary score evaluates other questions based on the original responses of the legal agent. If the ground truth is legal provisions, we measure the score by a matching method: 10 if the legal provisions are cited by the legal agent, and 0 if not. For ground truths that involve legal phrases, we leverage GPT-4o to generate a score ranging from 0 to 10. The evaluation prompt is shown in Figure 32.
@@ B IN N BIN B IN N BIN F OR D OC F OR D OC P FS J UD R EA L AW P FS C RI J UD R EA L AW
Table 5: performance among LLM-driven legal agents on J1-E NVS driven by Qwen3-Instruct-32B.
@@ D.2 Level-II
We evaluate the legal document drafted by the legal agent from two perspectives.
(1) Format-following score measures the correspondence of the generated document format with the given template. We break down the format into a set of labels $L=[L_{1},L_{2},...,L_{k}]$, which are organized in a specific sequential order $S$. As shown in Figure 29, the complaint format consists of labels of Plaintiff, Defendant, Claims, Facts and Legal Grounds, and evidence and Sources, Names and Addresses of Witness, following a specific sequential order. As for the defence format, check Figure 30, which includes Defendant, defence arguments, case id, parties and cause of action, evidence and Sources, names and addresses of witnesses. Through matching, we check whether the generated sequential order $S_{agent}$ follows the given order $S$, and score the accuracy of labels by comparing the generated labels $L_{agent}$ with $L$. The final result is calculated with the following equation.
@@
(2) Document score evaluates the quality of obtained information. To be specific, personal information, including the plaintiff’s information and the defendant’s information (i.e., name, ethnicity, address, birth date, gender), is measured through exact match(1 or 0). As for the other components(i.e., Claims, evidence, Statement of Defence, Case Details), they are evaluated with prompts in Figure 33.
@@ D.3 Level-III
(1) Procedural-following score provides assessments of the completeness of the procedural stages, each with several mandatory actions. Only when all mandatory actions are executed is a stage considered complete. Specifically, the civil Court is divided into five stages, i.e., court preparation, court investigation, court debate, court mitigation, and court decision. As illustrated in Figure 34 and Figure 35, we evaluate the court preparation stage, court investigation stage, court debate stage, and the court mitigation stage. On the other hand, as illustrated in Figure 36, Criminal Court is divided into three stages: court preparation, court investigation, and court decision. In addition, we notice that some legal agents are prone to unilateral speech, with all the stages executed within one turn. This performance is deemed a failure in our settings. Through matching, we calculate the completeness of the procedural stages and mandatory actions, and use their average as the $PFS$ score.
(2) Judgment score evaluates the quality of the legal agent’s judgment, which encompasses two parts: the inference part and the judgment part. The inference part contains the court’s inference for the judgments as well as the applied laws. We compare the laws applied by the legal agent with the ground truth to determine the correctness, and leverage GPT-4o to evaluate the inference process. The judgment part in the Civil Court includes the agent’s ruling over the case, which is evaluated with GPT-4o. The evaluation prompt could be seen in Figure 33. As for the Criminal Court, the judgment part is divided into three parts, where the criminal charge is scored through matching, and the accuracy of predicted fines and sentences is calculated with log deviation with the following equation. The average of the fine deviation and the sentence deviation is calculated as the final $VER$.
@@
Where: $v$ is the logarithmic deviation; $l_{p}$ is the agent predicted result; and $l_{a}$ is the ground truth. Higher $v$ indicates greater deviation and thus worse performance of the legal agent.
@@
For adapted LLMs, we speed up inference using vllm Kwon et al. [2023]. Greedy decoding was used across the evaluations. For models smaller than 32B, we run evaluations using 8 * RTX 4090 GPUs with 24GB memory each. For models larger than 32B, we run evaluations using 8 * A100 GPUs with 80GB memory. For some metrics calculating in the evaluation, we use GPT-4o 333gpt-40-2024-11-20. To prevent infinite interaction loops caused by limited model capabilities, we set maximum interaction rounds for each scenario: 15 for knowledge 1 questioning, 10 for legal consultation, 20 for complaint drafting, 15 for defence drafting, 50 for civil, and 35 for criminal court.
@@ F.1 Performance among Various Legal Agents
  • Complaint & Defence Drafting. Table 6 and Table 7 report additional DOC results in Complaint Drafting and Defence Drafting. Both general and legal-specific models perform relatively well on five sub-metrics of DOC, indicating powerful information collection ability through the multi-turn conversation. However, agents driven by legal-specific or smaller models, such as LawLLM, encounter difficulty in accomplishing the assigned tasks, requiring further training.
  • Civil Court. Table 8 reports additional PFS results in the Civil Court. GPT-4o and Claude-3.7 demonstrate remarkable performance in STA and ACT, indicating strong procedural-following ability of these models, while legal-specific models, the inference model, and smaller models are unable to follow civil court procedures.
  • Criminal Court. Table 9 reports additional PFS and DOC results in the Criminal Court. GPT-4o, Claude-3.7, and Deepseek-v3 score 100 in UNI. In contrast, legal-specific models fail to navigate the criminal court, resulting lower UNI and poorer performance.
@@ F.2 Analysis of Dialogue Turn Constraints on Simulated Courts Performance
We study the impact of maximum dialogue turns constraint on the performance of the legal agent in civil court and criminal court driven by Qwen3-Instruct-32B. For civil court, three groups are introduced, where the maximum dialogue turns are set to 35, 40, and 50. For criminal court, we examine legal agent performance under the maximum turn constraint of 30 and 35 turns. As illustrated in Figure 8, the performance of GPT-4o in civil court demonstrates a positive correlation with the increase in dialogue turn constraint, while Deepseek-v3 and Gemma3-Instruct-12B exhibit a decreasing trend with the increase in maximum dialogue turns. The gap in performance could be attributed to the difference in long-range memory, with GPT-4o effectively retaining context and executing procedural operations in detail with prolonged dialogue, while Deepseek-v3 and Gemma3-Instruct-12B suffer from the retention of early dialogue information, subsequently leading to diminished reasoning and judgment prediction performance. For the criminal court, as shown in Figure 9, across all evaluated agents, their performance remains largely stable with the increasing dialogue constraint, as most of cases are concluded within 30 turns. The results demonstrate the value of long-range memory for legal agent performance in complex and multi-turn environment settings, which highlights the necessity for future research on a more robust memory mechanism for legal agents.
@@
Table 10 presents the average number of interactions observed across the six environments. Conversely, agents that fail to complete the task tend to have few turns or directly reach the maximum turn limit. For instance, ChatGLM, InternLM3, and LawLLM achieve only 5.54, 1.05, and 2.45 turns on average at C I . The results indicate that weak interactive abilities prevent the models from sustaining focused, purposeful dialogues, often leading to off-topic, incomplete, or procedurally flawed interactions.
@@
  • Across different models. We select three representative models(i.e., GPT-4o, Deepseek-v3 and Gemma3-Instruct-12B), and re-run each model three times under the same settings of the main experiment in both the CD and DD scenarios. As shown in Table 13, performance of these models remains consistent across three runs, demonstrating the robustness of our framework.